An alkane has a molecular mass of 72. Draw its possible chain isomers.

Grade:12

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Deva

First of all, we will derive the molecular formula.

Molecular formula of alkane is CnH2n+2

As molecular mass = 72,

12n + 2n + 2 = 72

n = 5

The alkane is C5H12. The possible chain isomers are

 

chain isomers of C5H12

and

 

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

lanoaha brady
10 years ago
you are a dumbass this question doesn`t even make sense. I don`t understand why you would even try to evaluate this question????????
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Molecular formula of alkane is CnH2n+2
As molecular mass = 72,
12 x n + 1 x (2n + 2) = 72
or, 14n + 2 = 72
n = 5
The alkane is C5H12.
Hence, the possible chain isomers are the chain isomers of C5H12
 
Thanks and regards,
Kushagra